OpenStudy (anonymous):

Please help!! Simplify the expression by using a double angle formula cos^2(40) - sin^2(40)

OpenStudy (anonymous):

this can be written as cos(80)

OpenStudy (anonymous):

did u get it @ariabstr

OpenStudy (anonymous):

I dont know how to answer it now

OpenStudy (anonymous):

would it be \[\sin \frac{ \pi }{ 18 } - \cos \frac{ \pi }{ 18 }\]

OpenStudy (anonymous):

@KissMyAxe

OpenStudy (er.mohd.amir):

cos^2 x-sin^2 x=cos 2x

OpenStudy (anonymous):

@Er.Mohd.AMIR has explained it!

OpenStudy (er.mohd.amir):

its formula